# 4Université de Strasbourg, CNRS UMR 7550, Observatoire astronomique de Strasbourg, 67000 Strasbourg, France 5DIO, Observatoire de Paris, CNRS, PSL, 75014, Paris, France 6IDRIS, CNRS, F-91403 Orsay, France

### v1

@@ Appendix D Bootstrap from a degraded selection function
selecting 984 sources as detectable, which represents approximately 70% of the sources used for our original BASE training. Selecting only brighter sources should accelerate model training convergence, but is likely to result in a model unable to detect fainter sources. As we did for our reference training pipeline, we apply two successive bootstrap training steps, resulting in a total of three trained models, B-BASE, B-BT1, and B-BT2. We summarize the results in Table 6.
The B-BASE model catalog scores 6% below our BASE model with a 4% drop in the number of matches and a lower average source score, which is consistent with a reduced training parameter-space coverage. After a single bootstrap step, the B-BT1 model already achieves a similar score to our BASE model. The second bootstrap further improves the score by placing B-BT2 in between our BASE and BT1 models, confirming our hypothesis. Achieving higher scores with additional bootstrap steps would likely require some adjustments in our bootstrap strategy. As detailed in Sect. 3.4, all new sources are flagged as difficult instead of confident targets, which gives more weight to sources from the original selection function compared to the new ones, which is an issue if the starting selection function is too conservative. When generalizing this approach to observational data, we would recommend refining the process to remove the difficult flag of some targets based on their predicted objectness score.
Table 6: . SDC2 scores for models trained from a degraded selection function.
@@
Appendix E Ancillary models for transfer learning
In this paper, we followed the original challenge conditions as closely as possible to produce results that can be compared with other teams. Here, we aim to take a step forward in generalizing our method to observed hyperspectral datasets by distributing models pre-trained on the whole MAIN cube. For this purpose, we modify our training pipeline to generate examples from the MAIN cube at each iteration. This results in a much larger training sample with many detectable sources to train from than what we had with the LDEV cube. This would technically allow us to constrain a more complex and wider backbone and to relax some hyperparameters. However, it would require a new tedious exploration of the associated parameter space. In addition, using the small LDEV catalog as a validation dataset makes it difficult to compare the performances of different setups accurately. Here, we instead use the exact setup from our LDEV cube training, for which we have confirmed the high level of detection performance. We use the same starting selection function as our BASE training, and perform two bootstrap steps, resulting in three MAIN cube models: MC-BASE, MC-BT1, and MC-BT2.
The inference pipeline works the same way as before, but is applied to the smaller LDEV cube instead to monitor for blatant overtraining. The objectness selection threshold is optimized against the LDEV score. We summarize LDEV scorer-based results in Table 7. Due to the small number of LDEV sources, the scores we obtain are difficult to compare and likely sensitive to retraining variability. Still, we observe a clear 5% score improvement between MC-BASE and MC-BT1, while the additional score increase provided by MC-BT2 is negligible. We distribute these three models publicly, along with all other models presented in the paper, at xx.xxxx/zenodo.xxxxxxxx555 archive at https://share.obspm.fr/s/swyCT7BgEGjtZK3. These MAIN-cube models should be used preferentially when transferring SDC2 pre-trained models to new datasets, as they have been exposed to greater example diversity. They are also the models used in the interactive SDC2 inference notebook provided in the CIANNA git repository, as it is practically easier to do the inference over the small LDEV cube.
Table 7: . LDEV SDC2 scores for models trained on the MAIN cube.

### v2

@@ Appendix E Bootstrap from a degraded selection function
selecting 984 sources as detectable, which represents approximately 70% of the sources used for our original BASE training. Selecting only brighter sources should accelerate model training convergence, but it is likely to result in a model that cannot detect fainter sources. We then applied two successive bootstrap training steps, yielding three trained models: B-BASE, B-BT1, and B-BT2. We summarize the results in Table 7.
The B-BASE model catalog scores 6% below our BASE model with a 4% drop in matches and a lower average source score, which is consistent with a reduced training parameter-space coverage. After a single bootstrap step, the B-BT1 model already achieves a score similar to that of our BASE model. The second bootstrap further improves the score by placing B-BT2 in between our BASE and BT1 models, confirming our hypothesis. Here, we followed the approach described in Sect. 3.5 to add new sources to the training sample, thereby limiting the gains from the degraded selection function. In particular, all new sources are flagged as difficult rather than confident targets, giving more weight to sources from the original selection function than to the new ones, which is problematic if the starting selection function is too conservative. To address this issue, the selection process after a bootstrap step could be modified to include criteria for adding a source to the confident target list. For example, if a source is confidently detected over multiple consecutive bootstrap steps, it can be considered easy enough to detect. Also, in our current approach, we suppose that a source from the initial selection function is necessarily detectable, which might not be true. We already stated that sources originally selected that do not match the previous model’s predictions could be flagged as difficult, and if they are never detected across multiple bootstrap steps, they should likely be removed. This would be particularly useful for applications involving artifacts that can be selected by an initial function based on brightness but that do not match the morphology of typical sources from the training dataset.
Table 7: SDC2 scores for models trained from a degraded selection function.
@@
Appendix F Ancillary models for transfer learning
In this paper, we followed the original challenge conditions as closely as possible to produce results that can be compared with other teams. Here, we aim to take a step toward generalizing our method to observed hyperspectral datasets by distributing models pre-trained on the MAIN cube. For this purpose, we modified our training pipeline to generate examples from the MAIN cube at each iteration, resulting in a much larger training sample than with the LDEV cube. This change would technically allow one to constrain a more complex backbone and to relax some hyperparameters. However, it would require a tedious new exploration of the associated parameter space. In addition, using the small LDEV catalog as a validation dataset makes it difficult to accurately compare the performance of different setups. For these reasons, we chose to keep the exact setup from our LDEV cube training, which had already demonstrated high detection performance. We used the same starting selection function as for our BASE training and performed two bootstrap steps, producing three MAIN cube models: MC-BASE, MC-BT1, and MC-BT2.
The inference pipeline works the same way as before, and periodic predictions over the LDEV cube were used to monitor for blatant overtraining. The objectness selection threshold was optimized against the LDEV score for each model. We summarize LDEV scorer-based results in Table 8. Given the small number of LDEV sources, the scores we obtain are difficult to compare and likely sensitive to retraining variability. Still, we observe a clear 5% score improvement between MC-BASE and MC-BT1, while the increase from MC-BT1 to MC-BT2 is within typical retraining variability. We distribute these three models publicly, along with all other models presented in the paper, at 10.5281/zenodo.18403011. These MAIN-cube models should be used preferentially when transferring SDC2 pre-trained models to new datasets, as they have been exposed to greater example diversity. They are also the models used in the interactive SDC2 inference notebook in the CIANNA GitHub repository, since inference on the small LDEV cube is faster and easier to fit within limited computing resources.
Table 8: LDEV SDC2 scores for models trained on the MAIN cube.